# Strong-field Gravitational Wave Lensing in the Kerr Background

### v1

@@ IV Conclusion
By propagating the scattered field to a distant observer and coherently combining it with the direct waveform, we quantify how strong-field wave-optics Kerr lensing modifies incident GW signals. The effects are most pronounced when the source and lens lie sufficiently close to each other, in which case the scattered contribution arrives within milliseconds of the direct signal, which is well inside the duration of typical LIGO event, and produces noticeable amplitude and phase modulations. To quantify this more concretely, we performed a mismatch analysis for certain source––lens––observer configurations. These results show that strong-field scattering effects could be detectable even with current detectors, and the improved sensitivity and higher SNR of future observatories will only strengthen this prospect. In particular, as already discussed in the Sec. I, percentage-level mismatches found for $\theta_{o}=30^{\circ}$ lie near the detectability threshold of the current ground-based detectors and may be observable in sufficiently loud events. Looking ahead, next-generation detectors such as ET, CE, and LISA will achieve SNRs in the hundreds to thousands, reducing the detectable mismatch threshold to $10^{-4}-10^{-6}$. In these regimes, the strong-field Kerr-lensing signatures predicted here become decisively measurable, opening a realistic pathway to probing strong-field wave-optical effects in astrophysical environments.
@@ Acknowledgements.
We thank Uddeepta Deka for helping us perform the mismatch analysis and providing valuable feedback. We also thank Chris Kavanagh and Ajith Parameswaran for useful comments during the work. We extend our gratitude to the Astrophysical Relativity group at ICTS, for many engaging discussions and useful feedback. We acknowledge the use of IUCAA LDG cluster Sarathi for the computational work. The work utilizes Mathematica [Mathematica] and the following software packages: Black Hole Perturbation Toolkit [BHPToolkit], Cython [behnel2011cython], FLINT [FLINT], Jupyter notebook [Kluyver2016jupyter], LALSuite [2020ascl.soft12021L], Matplotlib [Hunter:2007], NumPy [Harris:2020xlr], PyCBC [pycbc_github], and SciPy [Virtanen:2019joe]. The research of M.V.S.S is supported by the National Post-Doctoral Fellowship (PDF/2025/004764), ANRF, Government of India.
@@ Appendix B Derivation of the phase formula
$\eta^{\rm P}_{lm}(\omega^{\prime})e^{2i\delta^{\rm P}_{lm}(\omega^{\prime})}=(-1)^{l+1}\frac{\Re(C)+12iM\omega^{\prime}~{\rm P}(-1)^{l}}{16\omega^{\prime 4}}\frac{B^{(\text{refl})}_{lm\omega^{\prime}}}{B^{(\text{inc})}_{lm\omega^{\prime}}}$ (54)
@@ Appendix D Computing correction to waveform in frequency domain
In the main text, we showed how a monochromatic-plane wave is scattered by a Kerr BH lens. Here, we generalize this to the case where the plane wave is instead a superposition of frequencies, as expected from realistic GW signals that evolve in time. It is simplest to work in the frequency domain. We again restrict to the case where the lens experiences purely positive helicity GW radiation, for simplicity. It is straightforward to generalize this setup.

### v2

@@ IV Conclusion
By propagating the scattered field to a distant observer and coherently combining it with the direct waveform, we quantify how strong-field wave-optics Kerr lensing modifies incident GW signals. The effects are most pronounced when the source and lens lie sufficiently close to each other, in which case the scattered contribution arrives within milliseconds of the direct signal, which is well inside the duration of typical LIGO event, and produces noticeable amplitude and phase modulations. To quantify this more concretely, we performed a mismatch analysis for a class of source–lens–observer configurations. These results show that strong-field scattering effects could be detectable even with current detectors for a subset of considered configurations, and the improved sensitivity and higher SNR of future observatories will only strengthen this prospect. In particular, as already discussed in the Sec. I, percentage-level mismatches found for $\theta_{o}=30^{\circ}$ lie near the detectability threshold of the current ground-based detectors and may be observable in sufficiently loud events. Looking ahead, next-generation detectors such as ET, CE, and LISA will achieve SNRs in the hundreds to thousands, reducing the detectable mismatch threshold to $10^{-4}-10^{-6}$. In these regimes, the strong-field Kerr-lensing signatures predicted here become decisively measurable, opening a realistic pathway to probing strong-field wave-optical effects in astrophysical environments.
@@ Acknowledgements.
We thank Uddeepta Deka for helping us perform the mismatch analysis and providing valuable feedback. We thank Chris Kavanagh and Ajith Parameswaran for useful comments at many phases of the work. We extend our gratitude to Luka Vujeva and Juno Chan for carefully reading an earlier draft and providing valuable feedback. We are also thankful to the Astrophysical Relativity group at ICTS, for many engaging discussions and useful feedback. We acknowledge the use of IUCAA LDG cluster Sarathi for the computational work. The work utilizes Mathematica [Mathematica] and the following software packages: Black Hole Perturbation Toolkit [BHPToolkit], Cython [behnel2011cython], FLINT [FLINT], Jupyter notebook [Kluyver2016jupyter], LALSuite [2020ascl.soft12021L], Matplotlib [Hunter:2007], NumPy [Harris:2020xlr], PyCBC [pycbc_github], and SciPy [Virtanen:2019joe]. The research of M.V.S.S is supported by the National Post-Doctoral Fellowship (PDF/2025/004764), ANRF, Government of India. The research of RG and AM is supported by the Department of Atomic Energy, Government of India, under Project No. RTI4001.
@@ Appendix B Derivation of the phase formula
$\eta^{\rm P}_{lm}e^{2i\delta^{\rm P}_{lm}}(\omega^{\prime})=(-1)^{l+1}\frac{\Re(C)+12iM\omega^{\prime}~{\rm P}(-1)^{l}}{16\omega^{\prime 4}}\frac{B^{(\text{refl})}_{lm\omega^{\prime}}}{B^{(\text{inc})}_{lm\omega^{\prime}}}$ (57)
@@ Appendix D Computing correction to waveform in frequency domain
In the main text, we have already shown (via SFSF) how a monochromatic-plane wave is affected by a Kerr BH lens. Here, we generalize this to the case where the plane wave is instead a superposition of frequencies, as expected from realistic GW signals that evolve in time. It is simplest to work in the frequency domain. We again restrict to the case where the lens experiences purely positive helicity GW radiation, for simplicity. It is straightforward to generalize this, following the discussion below Eq. (18), in cases where the incident wave is a linear combination of both helicities.